### v1

# FonTS: Text Rendering with Typography and Style Controls

Abstract
Visual text images are prevalent in various applications, requiring careful font selection and typographic choices. Recent advances in Diffusion Transformer (DiT)-based text-to-image (T2I) models show promise in automating these processes. However, these methods still face challenges such as inconsistent fonts, style variation, and limited fine-grained control, particularly at the word level. This paper proposes a two-stage DiT-based pipeline to address these issues by enhancing controllability over typography and style in text rendering. We introduce Typography Control (TC) finetuning, an efficient parameter fine-tuning method, and enclosing typography control tokens (ETC-tokens), which enable precise word-level application of typographic features. To further enhance style control, we present a Style Control Adapter (SCA) that injects style information through image inputs independent of text prompts. Through comprehensive experiments, we demonstrate the effectiveness of our approach in achieving superior word-level typographic control, font consistency, and style consistency in Basic and Artistic Text Rendering (BTR and ATR) tasks. Our results mark a significant advancement in the precision and adaptability of T2I models, presenting new possibilities for creative applications and design-oriented tasks.
{strip}[Uncaptioned image]
[…]
1 Introduction
Visual text images are ubiquitous in daily life and hold significant commercial value in advertising, branding, and marketing [3, 9]. However, the design process for visual text is complex and time-consuming. Designers must carefully select appropriate fonts, use typographic elements like italics, and create artistic styles that are aesthetically pleasing and coherent. Recent advances in diffusion models [31, 27] demonstrate promising potential for creating visual contents in design, thereby attracting substantial attention. Concurrently, real-world applications raise increasing demands for controllability over the generated content.
Previous efforts have mainly focused on improving control over the accuracy of scene text rendering [44, 38, 8, 9]. With the development of DiT-based T2I models, e.g. SD3 [12] and Flux .1 [1], the accuracy of text rendering has seen significant improvements. Beyond content accuracy, Glyph-ByT5 [20] introduced a new text encoder pre-trained using contrastive learning, enabling various font types in text rendering. Textdiffuser-2 [9] trained both two language models and the whole diffusion model to acquire layout planing capabilities. While these methods [20, 9] have implemented control at the paragraph-level (more than 10 words), they have not yet realized word-level control. Moreover, prior text rendering methods often overlook the artistic aspects of text rendering [9]. Recent DiT models [12, 1] have demonstrated promising capabilities in artistic text rendering, yet they face challenges such as semantic confusion and style inconsistency.
To expand the boundaries of existing methods and provide enhanced control on image generation, this paper identifies three essential requirements that text rendering methods: 1) control of fonts and word-level attributes in Basic Text Rendering (BTR); 2) consistency in style control in Artistic Text Rendering (ATR); 3) preservation of Scene Text Rendering (STR) capabilities without negative impact.
To this end, we propose a two-stage DiT-based pipeline for text rendering with typography and style controls. To achieve typography control, we introduce Typography Control (TC)-finetuning, an efficient parameter fine-tuning method, alongside enclosing typography control tokens (ETC-tokens) and paired typography control dataset (TC-dataset). Our results demonstrate that the model not only learns typographic elements but also applies specific typographic features at precise word locations. To ensure content accuracy while enhancing style control, we introduces a style control adapter (SCA) that injects style information via images, independent of the text content in the prompt. For training the SCA, we collected a dataset of approximately 600k image-text pairs with high aesthetic scores.
We validate the effectiveness of the proposed methods. First, we demonstrate that the learned ETC-tokens can generate text images with the desired word-level typographic attributes, through GPT-4o [26] and manual verification. Next, we assess font consistency in BTR and style consistency in ATR by conducting user studies. These evaluations show that our method outperforms various baseline methods in terms of font consistency and word-level controllability for BTR, as well as style consistency for ATR.
In summary, our contributions are as follows:
  • We present a DiT-based pipeline that enhances the controllability of typography and style in text rendering while preserving the capabilities of scene text rendering.
  • We propose a parameter-efficient technique that enables DiT-based T2I models to learn and apply word-level typographic attributes. The use of ETC-tokens allows model to learn specific attributes and the precise locations.
  • Comprehensive experiments validate the effectiveness of our method via superior performance in typography and style control compared to existing baselines. This is the first work to achieve word-level control in text rendering and provides a valuable reference for future research.
[…]
2 Related Work
Scene Text Rendering. Despite progress in diffusion models [31, 27], high-quality scene text rendering remains a challenge. To address this, one line of research [44, 9, 8, 37] focuses on explicitly controlling the position and content of the text being rendered. For instance, methods such as GlyphControl [44] and AnyText [37] leverage glyph images containing multiple text lines as priors to guide diffusion models to produce accurate accurate text, relying on ControlNet [46]. Similarly, TextDiffuser [9, 8] employs character-level segmentation masks as control conditions in scene text rendering. Another line of works [20, 21] fine-tune the character-aware ByT5 encoder [19] using paired glyph-text datasets, improving the ability to render accurate text in various real-world images.
While these methods improve text content accuracy in scene text rendering, they have led to a trade-off in model performance on stylization, limiting text style diversity and complicating the rendering of artistic text images [8].
Artistic Text Rendering. Early research focused on font creation by transferring textures from existing characters, employing stroke-based methods [5, 33] and patch-based techniques [40, 41, 42]. The advent of Generative Adversarial Networks (GANs) improved the synthesis of new glyphs from limited examples and the separation of style and content in font design, enhancing flexibility and realism [2, 15, 14, 43, 22, 35]. Innovations with diffusion models [36, 39, 25] have enabled diverse text image stylization and semantic typography, resulting in visually appealing designs that retain readability. However, despite recent DiT models [12, 1] showing quite promise in artistic text rendering, they still struggle with semantic confusion and style inconsistency.
Our work differs from previous research: 1) Rather than stylizing an existing text glyph image, we generate artistic text images based on the text rendering capabilities of the DiT model (Flux [1]). 2) We trained style adapters on large-scale image-text pairs to handle various style conditions, including artistic text images and arbitrary style images.
Controllable Text Rendering and Image Generation. Controllable text rendering. Glyph-ByT5 [20] introduces a text encoder trained through contrastive learning, enabling paragraph-level (more than ten words) control over font types. Textdiffuser-2 [9] trains two language models with the full U-Net backbone to achieve layout planning capabilities for text rendering. While these methods have successfully enabled paragraph-level control, challenges remain in achieving fine-grained word-level control.
Text-based controllable generation. Text-based controllable generation methods customize image outputs by fine-tuning diffusion models using user-provided examples. [13, 32, 16]. These approaches introduce modifier tokens to guide the generation process. ColorPeel [6] further enhances this by constructing datasets of color-shape pairs and corresponding color-prompt tokens, enabling the generation of images with precise target colors. Unlike methods that use a single modifier token to represent an object [16, 13] or a color [6], we propose using enclosing modifier tokens. Each enclosing token consists of a starting and an ending token to represent typographic attributes. This approach allows the model to not only learn the attribute but also understand its precise application location, such as a specific word.
Image-based controllable generation. Methods for image-based control, such as UniControl [28] retrain T2I models from scratch, are computationally expensive [46]. A efficient alternative introduces trainable modules to existing architectures [24, 45, 46, 47]. These adapter-based methods enable structural control [46, 24] and style control from images [45, 7]. Most of them have been implemented on U-Net backbones, with limited exploration of DiT-based T2I models. Additionally, controllable generation based on multi-modal conditions remains underexplored.
Our work extends existing controllable generation methods to DiT-based T2I models, enabling fine-grained, word-level typographic control. We also integrate both text-based control and image-based control into the generation process, broadening the applicability in generated contents.
3 Approach
Our proposed parameter-efficient fine-tuning method with enclosing typography control tokens (ETC-tokens), shown in Figure 2 (a). In addition, we detail the training of style control adapters for DiT blocks of the transformer backbone, depicted in Figure 2 (b).
3.1 Typography Control Learning
Preliminaries of Rectified Flow DiT. To avoid the computationally expensive process of ordinary differential equation (ODE), diffusion transformers such as [12, 1] directly regress a vector field $u_{t}$ that generates a probability path between noise distribution $p_{1}$ and data distribution $p_{0}$. To construct such a vector field $u_{t}$, [12] consider a forward process that corresponds to a probability path $p_{t}$ transitioning from $p_{0}$ to $p_{1}=\mathcal{N}(0,1)$. This can be represented as $z_{t}=a_{t}x_{0}+b_{t}\epsilon\quad\text{, where}\;\epsilon\sim\mathcal{N}(0,I)$. With the conditions $a_{0}=1,b_{0}=0,a_{1}=0$ and $b_{1}=1$, the marginals $p_{t}(z_{t})=\mathbb{E}_{\epsilon\sim\mathcal{N}(0,I)}p_{t}(z_{t}|\epsilon)\;$ align with data and noise distribution. Referring to [18, 12], the marginal vector field $u_{t}$ can generate the marginal probability paths $p_{t}$, using the conditional vector fields as follows:
$u_{t}(z)=\mathbb{E}_{\epsilon\sim\mathcal{N}(0,I)}u_{t}(z|\epsilon)\frac{p_{t}(z|\epsilon)}{p_{t}(z)},$ (1)
The conditional flow matching objective is formulated as:
${L_{CFM}}=\mathbb{E}_{t,p_{t}(z|\epsilon),p(\epsilon)}||v_{\Theta}(z,t)-u_{t}(z|\epsilon)||_{2}^{2}\;,$ (2)
where the conditional vector fields $u_{t}(z|\epsilon)$ provides a tractable and equivalent objective.
Figure 3: Comparative weight changes in the transformer backbone during full parameter fine-tuning. (a) shows that the MM-DiT experiences double the weight changes compared to the Single-DiT. (b) indicates that the Txt-Attn also shows the double weight changes relative to other components within the MM-DiT.
Typography Control Fine-tuning. While previous studies have shown that fine-tuning specific U-Net components can generate specific objects and colors through learned prompts [6, 16, 13], approaches are not suitable for our task due to architectural differences between DiT and U-Net. Following [16, 17], we analyzed parameter changes in the fine-tuned transformer backbone on the target dataset for 100k steps using the loss ${L_{CFM}}$ in Eq. 2. The change in parameters for layer $l$ is calculated as $\Delta_{l}=||\theta_{l}^{\prime}-\theta_{l}||/||\theta_{l}||,$, where $\theta_{l}^{\prime}$ and $\theta_{l}$ are the fine-tuned and pretrained model parameters, respectively. The total change across all layers is: $\Delta_{sum}=\sum_{l=0}^{n}mean(\Delta_{l})$. These parameters are derived from two types of layers: (1) MM-DiT blocks (merging text and image embeddings), and (2) Single-DiT blocks (processing merged embeddings from MM-DiT). In MM-DiT blocks, parameters are divided into three components: joint text attention (Txt-Attn), joint image attention (Img-Attn), and additional modules like multi-layer perception (MLP) and modulation blocks. Figure 3 shows that MM-DiTs have approximately double the weight change of Single-DiTs, with the Txt-Attn component showing nearly twice the change of other MM-DiT elements, despite it is only 5% of the total parameters in the transformer backbone.
Figure 4: Examples of TC-Dataset featuring two types of TC-Tokens. (a) illustrates the TC-token for various font types. (b) displays the ETC-token with word-level typographic attributes applied to a specific word, including bold, italic, and underline.
[…]
To fill the gap in high-quality datasets that combine text with typographic attributes, we created the TC-Dataset using typography control rendering (TC-Render). The pipeline leverages HTML rendering to produce images featuring typographic attributes like fonts and word-level styles such as bold, italic, and underline, as shown in Figure 4. Details of the dataset are available in the supplementary.
3.2 Style Control Adapters
Decoupled Joint Attention. The joint attention here refers to the attention in MM-DiT blocks of SD3 [12] and Flux [1]. Given the text features $c_{txt}$ and input of joint attention $z_{t}$, the output of joint attention ${z}^{\prime}$ can be defined as:
$\vspace{-1.5mm}\begin{split}{z}^{\prime}=\text{Attention}({Q},{K},{V})=\text{Softmax}(\frac{{Q}{K}^{\top}}{\sqrt{d}}){V},\\ \end{split}$ (3)
where ${Q}=z_{c}{W}_{q}$, ${K}=z_{c}{W}_{k}$, ${V}=z_{c}{W}_{v}$ are the query, key, and values matrices of the attention operation respectively, $z_{c}=concat({z_{t}},{c_{txt}})$,and ${W}_{q}$, ${W}_{k}$, ${W}_{v}$ are the weight matrices of the trainable layers.
In order to better decouple style and content, we additionally introduce a decoupled joint attention mechanism (DJA). Inspired by [45, 7, 24], we add DJA at the joint attention layers for text features $c_{txt}$ and image features $c_{img}$ are separate. To be specific, we add new joint attention layers in the original MM-DiT and Single-DiT blocks to insert image features. Given the image features $c_{img}$, the output of new joint attention ${z}^{\prime\prime}$ is as follows:
$\vspace{-1.5mm}\begin{split}{z}^{\prime\prime}=\text{Attention}({Q}^{\prime},{K}^{\prime},{V}^{\prime})=\text{Softmax}(\frac{{Q^{\prime}}{K^{\prime}}^{\top}}{\sqrt{d}}){V}^{\prime},\\ \end{split}$ (4)
where, ${Q}^{\prime}={z_{t}}{W}_{q}$, ${K}^{\prime}={c}_{img}{W}^{\prime}_{k}$ and ${V}^{\prime}={c}_{img}{W}^{\prime}_{v}$ are the query, key, and values matrices from the image features. ${W}^{\prime}_{k}$ and ${W}^{\prime}_{v}$ are the corresponding weight matrices. Consequently, we only need to add two parameters ${W}^{\prime}_{k}$, ${W}^{\prime}_{v}$ for each decoupled joint attention layer. Then, we simply add the output of image cross-attention to the output of text cross-attention. Hence, the final formulation of the decoupled cross-attention is defined as follows:
$\begin{split}{z}^{new}=\text{Softmax}(\frac{{Q}{K}^{\top}}{\sqrt{d}}){V}+\lambda*\text{Softmax}(\frac{{Q^{\prime}}{K^{\prime}}^{\top}}{\sqrt{d}}){V}^{\prime},\\ \end{split}$ (5)
where ${Q}={z_{c}}{W}_{q},{K}={z_{c}}{W}_{k},{V}={z_{c}}{W}_{v},{Q}^{\prime}={z_{t}}{W}_{q},{K}^{\prime}={c}_{img}{W}^{\prime}_{k},{V}^{\prime}={c}_{img}{W}^{\prime}_{v}$ , and $\lambda$ represents the influence of $c_{img}$. We freeze the base model, only the ${W}^{\prime}_{k}$ and ${W}^{\prime}_{v}$ are trainable.
Methods / $\frac{\textbf{Task}}{\textbf{Metrics}}$ Basic Text Rendering Artistic Text Rendering
[…]
Glyph-ByT5 [20] 96.36 - 32.73 - -
TextDiffuser-2 [9] 42.86 - 1.81 - -
SD3-medium [12] 48.05 - 0.91 24.16 13.64
Flux.1-dev [1] 66.49 - 0.91 48.71 17.42
Ours 82.85 55.00 63.64 61.78 74.43
Table 1: Comparison of quantitative results and user study outcomes. The best results are indicated in bold formats.
[…]
Flux.1-dev [1] 24.17 29.93
Ours 53.57 31.66
Table 2: Quantitative results for scene text rendering.
Figure 5: Qualitative results on the font consistency and word-level controls in basic text rendering compared with baselines.
Figure 6: Qualitative results on the artistic text rendering task compared with Midjourney [23], Flux [1] and SD3 [12].
[…]
Txt-Attn(Ours) 5.10% 81.43 55.00
Txt+Img-Attn 9.40% 77.92 31.00
Table 3: Ablation studies on the selection of parameters during TC-finetuning. Assessing the basic text rendering capabilities with different trained components of the DiT backbone.
[…]
Ours w/ TC-FT 54.29 63.75
Ours w/o TC-FT 41.90 16.50
Table 4: Ablation studies examining the impact of TC-finetuning (TC-FT) on artistic text rendering.
Figure 7: Qualitative results on the style consistency in artistic text rendering compared with Flux [1] and SD3 [12].
4 Experiments
Text Rendering Benchmark. To assess the text rendering capabilities with word-level typography and style controls, we extend the existing scene text rendering benchmark [8] by introducing new benchmarks for basic text and artistic text rendering. BTR-bench. To evaluate word-level typography controls in basic text rendering, we introduce the Basic Text Rendering benchmark (BTR-bench). BTR-bench includes 100 prompts of different fonts and typographic attributes. For each text prompt, typographic attributes are randomly applied to three positions within the text to assess the model’s ability to render specific typographic attributes on individual words, while font attributes are applied to the entire text in the image. ATR-bench. To evaluate artistic text rendering, we introduce the Artistic Text Rendering benchmark (ATR-bench). Based on the single-letter and multi-letter classification in [36], we categorize the content into single-word and multi-word groups. Drawing on the style prompts from the GenerativeFont benchmark [25], we generate artistic individual letters and words using Flux [1]. These generated artistic letters and words are used for single-word and multi-word text rendering, respectively.
Style Control Dataset. Style control training consists of two phases, each utilizing different datasets. The first phase involves pretraining with a dataset of general image-text pairs. The second phase is fine-tuning, using a dataset that includes artistic text images and paired descriptions. For pretraining, we assembled a dataset called SC-general, which includes approximately 580k general image-text pairs with high aesthetic scores. These images were sourced from open-source datasets [11, 34]. For the fine-tuning phase, we created the SC-artext dataset. We compile a list of style descriptions and a list of words. Combining these lists generated various prompts for artistic text images, which were then used as input for Flux.1-dev [3], resulting in approximately 20k high-quality images. To ensure the images matched the original text content, we used shareGPT4v [10] to regenerate captions. Details of these two datasets can be found in the supplementary material.
Implementation Details. We use Flux.1-dev [1] as the base model. In TC-finetuning, we omit the text condition from the CLIP text encoder, as the pooled text embedding is not fed into the joint attention of the DiT backbone. For training the style control adapters, we select CLIP [29] as the image encoder since its visual embeddings are not sensitive to text [20], ensuring minimal impact on the content of the text while modifying the style.
Training details. For TC-finetuning, we fine-tune the base model for 40k steps using the TC-Dataset, incorporating a regularization prefix (‘sks’) in the text prompts. The batch size is set to 2. For the style control adapters, we train for 100k steps on the SC-general dataset and 15k steps on the SC-artext dataset, using a batch size of 4. All training is conducted on 8 A100 GPUs, with a learning rate of $1\times 10^{-5}$. To optimize memory usage, we backpropagate the loss every two steps and utilize DeepSpeed ZeRO [30].
4.1 Quantitative Results
For quantitative evaluation, we use the same OCR tool [4] as in [37] to detect text content in the generated images and calculate the OCR accuracy (OCR-Acc) across three distinct text rendering tasks. In the basic text rendering (BTR) task, existing OCR tools struggle to evaluate word-level typographic attribute accuracy (Word-Acc). Therefore, we use GPT4o and manual screening to assess and obtain the corresponding score. Font consistency (Font-Con) in BTR and style consistency (Style-Con) in artistic text rendering (ATR) are evaluated through user studies. Additionally, we compare our method to Flux.1 in the scene text rendering (STR) task (using the MARIO-bench [8]), evaluating OCR-Acc and CLIP scores (CLIP). Since Glyph-ByT5 and TextDiffuser-2 do not support the artistic text rendering defined in our study (we provide their qualitative results on the ATR-bench in the supplementary materials), the ATR column for these methods lacks quantitative results.
As shown in Table 1, our method outperformed the baselines in four out of five metrics while slightly below Glyph-ByT5 regarding OCR accuracy in BTR. This is reasonable since Glyph-ByT5 was trained on millions of text images, whereas our approach utilized a dataset of 50k basic text images, which is twenty times smaller.
Furthermore, as indicated in Table 2, our method significantly improves OCR-Acc and CLIP scores compared to Flux. Upon reviewing the image results, we found that Flux exhibits semantic confusion in the STR task (on MARIO-bench [8]), similar to its performance in ATR, which notably reduces its OCR accuracy.
User Studies. We conducted user studies with 22 participants to perceptually evaluate our results, comparing them against baseline methods. The evaluation included questions focused on two key aspects: font consistency (Font-Con) and style consistency (Style-Con). Font-Con was assessed through two subtypes: the first measured the consistency between the output image and the ground truth, while the second assessed font consistency across different outputs under identical font input conditions. Style-Con was evaluated similarly, with two subtypes covering the same aspects. The number of questions for each subtype is: 4, 2, 3, and 2, respectively. The score for each method was calculated based on the number of votes it received divided by the total number of votes. According to the results, our method achieved the best performance in the Font-Con and Style-Con categories, as shown in Table 1. Further details are provided in the supplementary materials.
4.2 Qualitative Results
Basic Text Rendering. We use a set of challenging prompt words for evaluation. For Flux.1 and Glyph-ByT5, the prompt example (for the leftmost images) is: “Blue Text: ‘Love knows no limits’ in Font: Josefin Sans, Add underline to ‘Love’, Background: pure yellow”. This prompt specifies the font applies a particular typographic attribute to a word and defines the text and background colors. As shown in Figure 5, Glyph-ByT5 achieves better font consistency than Flux but lacks word-level control. In contrast, our method ensures strong font consistency and enables word-level control, such as underline, bold, or italic.
Artistic Text Rendering. As TextDiffuser-2 and Glyph-ByT5 do not support artistic text rendering, We compare our method with SD3 and Flux. All qualitative results for our method are displayed with an image scale set to 0.9. We conduct solid experiments and have the following findings:
Semantic confusion. SD3 and Flux do not process image inputs, while ours is capable of handling both images and text prompts. To this end, we also report the results generated by Midjourney [23] using the same inputs as ours. As shown in Figure 6, in contrast to other approaches, our method effectively integrates style control without compromising the readability and integrity of the original text content. Meanwhile, it greatly mitigates the phenomenon of semantic confusion existing in all three comparison baselines.
Style consistency. To ensure that SD3 and Flux produce artistic text images with a consistent style, we set a fixed seed during generation and employed the same style prompt for all inputs, varying only the text content. Our method utilized the artistic single letter Flux generated with the same style prompt ( ‘A’ and ‘n’ in Figure 7), for style controls. The text prompt follows a simple format: Text: ‘content’ where content refers to the desired text for generation.
As depicted in Figure 7, results generated by SD3 have poor accuracy in some words, with errors in both letter transcription (e.g., ‘Banana’ becoming ‘BNANA’ in the second image of the first row) and case sensitivity (e.g., all uppercase letters in the sixth image of the first row). The middle row shows results from Flux, also revealing semantic confusion, notably for words with strong semantic like ‘parrot’, where the red parrot replaces the ‘p’ in ‘parrot’ as seen in the fifth image of the second row. The results in thelast row displays demonstrating superior text content accuracy and style consistency of our method compared to the baselines.
In addition, we find that the typography controls in BTR can be transferred to ATR and STR, as shown in  1 to a certain extent. It is reasonable that the degree of controllability will be affected by the given style image, particularly when it contains text. However, considering we only train the model using basic text images to learn those typography attributes, it shows considerable potential for generalization and domain adaption ability in the proposed method.
4.3 Ablation study
Ablation on TC-Finetuning. To assess the effectiveness of TC-finetuning, we performed ablation studies on the fine-tuned module, training two configurations: joint text attention (Txt-Attn) and joint text and image attention (Txt+Img-Attn). The results in Table 3 show the performance differences across these two settings for the BTR task. The data indicates that increasing the training parameters for the Img-Attn component actually reduces model performance on the BTR task, suggesting that fine-tuning only the Txt-Attn is a more effective approach. Meanwhile, we investigated the impact of TC-finetuning on artistic text rendering. By comparing the performance before and after finetuning, we observed a significant improvement in OCR accuracy, particularly for the multi-word subset, as shown in Table 5.
Ablation on ETC-Tokens. The ETC tokens are designed for assigning the words which need to be affected. To this end, assume we want to make the word ‘Find’ Bold of ‘Find your path’, we consider three cases: 1) no tokens: generate the results using language to express the intentions, i.e. using the prompt as ‘make the Find Bold in the content of Find your path’; 2) single token: following previous research [16, 6], we trained the model use single token and put the modifier token in front of the ‘Find’; 3) Ours. It can be found that the generated results using the first way brings great semantic confusion and the second way generated random results instead of accurately affecting the specific words. Details present in the supplementary materials.
4.4 Applications and Limitation
Applications. Artistic font design. Benefit from the robust style consistency, our approach is able to generate a variety of artistic letters with high consistency. Moreover, because the SCA is pre-trained on high-quality, large-scale data, the style control is not limited to artistic text images. Any style image can be used as a control input, as shown in Figure 8.
Stylization of scene text image. Since visual embeddings of CLIP [29] are not highly sensitive to text [20], using CLIP as the image encoder for style control adapters has minimal impact on the content of the text. Instead, it primarily affects the stylistic presentation. As shown in Figure 9, after injecting different style prompts, the style of text changes, but the content of text remains clearly identifiable.
Logo design. Scene text and artistic text images can also be seamlessly integrated. By using scene text image prompts alongside artistic text images for style control, our method achieves a smooth blend of the two, as shown in Figure 10. This allows for the creation of versatile logo designs that are suitable for a range of application scenarios.
Limitation. It is observed the language drift phenomenon exists in our method, as the same as [16, 32]. This effect becomes noticeable as the number of training steps increases. This is mainly because, in the TC-finetuning process, we did not use additional regularization datasets; instead, we applied a simple regularization prefix, ‘sks’, in the text prompts of the TC dataset. This way decreases the cost. As shown in Figure 11, although language drift is severe at 60k steps, leading to the separation of text and scene in the generated image, the results at 40k steps are acceptable.
Figure 8: The results of artistic letters with different styles.
Figure 9: The results of stylized scene text image with different images and image scales.
Figure 10: The logo design of stylized scene text image with artistic text images and different image scales.
Figure 11: Inference results of TC-finetuned models at different training steps. The example prompts are from MARIO-bench [8].
5 Conclusion and Future Work
This paper proposes a two-stage DiT-based pipeline for text rendering with typography and style controls. TC-finetuning with ETC-tokens enables the model to learn and apply word-level typographic attributes to specific words. Style Control Adapter allows for independent style control using images. Experimental results show that the proposed method outperforms baselines on font consistency and word-level control in basic text rendering (BTR) and style consistency in artistic text rendering (ATR). To our knowledge, this paper is the first to achieve word-level control in basic text rendering, offering a valuable reference for future research in both scene and artistic text rendering. In future work, we aim to refine the scalability of our method further to handle a broader range of typography controls and styles and explore its extension to multilingual text rendering.
Supplementary Material
{strip}[Uncaptioned image]
Figure 12: Results of our method: (a), (b) and (c) in basic text rendering, artistic text rendering, and scene text rendering, respectively.
Figure 13: Examples of TC-Dataset. (a) different word-level attributes, (b) examples featuring text and background color variations.
Figure 14: Examples of images in SC-general.
Figure 15: Examples of SC-artext. All these artistic text images are generated by Flux.1-dev [1].
Figure 16: Examples of Semantic Confusion in Flux.1-dev  [1]. The prompts for the right three images are from MARIO-bench [8].
Figure 17: Examples of Typographic Controls in Scene Text Rendering.
This supplementary material serves as a complement to the main paper, including demonstration of BTR, ATR and STR in Section A; details of the datasets used in Section B; details of user study in Section C; additional results presented in Section D; and further details regarding word accuracy (Word-Acc) in Section E.
Appendix A Demonstration of BTR, ATR and STR
This section provides additional information to complement Section 1 of the main paper, which outlines the scope of three text rendering tasks:
  • Basic Text Rendering (BTR) involves rendering simple text on a solid color background without any additional scene elements, as illustrated in Figure 12(a).
  • Artistic Text Rendering (ATR) features a minimalist background that highlights the artistic nature of the text itself, as seen in Figure 12(b).
  • Scene Text Rendering (STR) involves integrating text and scene elements in a way that shares contextual meaning and blends harmoniously, as depicted in Figure 12(c).
Appendix B Details of Datasets
This section complements Section 3 , 4 of the main paper, detailing the datasets we utilized in our study.
Typography Control Dataset (TC-Dataset). To address the lack of high-quality datasets that integrate text with word-level typographic attributes, we developed the TC-Dataset using typography control rendering (TC-Render). This process harnesses HTML rendering to generate images that display typographic features such as various fonts and word-level attributes, including bold, italic and underline. We initiated our process by extracting 625 text excerpts from novels. For each excerpt, we designed an HTML structure comprising sixteen images: one without typographic attributes and, in five different positions, applied three distinct typographic attributes (shown in Figure 13 (a)). Furthermore, we applied data augmentation techniques by randomly altering the text color and background (shown in Figure 13 (b)). Each HTML structure was rendered with one of five different fonts, resulting in approximately 50k text-image pairs set against solid color backgrounds.
Style Control Dataset (SC-Dataset).
SC-general. To train our style control adapters, we assembled the SC-general dataset, which includes approximately 580k general image-text pairs with high aesthetic scores. These pairs were sourced from open-source datasets [11, 34]. Figure 14 presents sample images, and Table 6 displays the corresponding paired texts.
SC-artext. For fine-tuning the style control adapters, we created the SC-artext dataset. We combined a list of 100 style descriptions with a list of 99 words, categorized into three character length groups: 1-15, 16-30, and 30-50. This combination produced a variety of prompts for artistic text images, which served as input for Flux.1-dev [1], yielding around 20k high-quality images. To ensure the images accurately reflected the original text content, we utilized shareGPT4v [10] to regenerate captions. Figure 15 shows sample images, and Table 6 presents the paired texts.
Appendix C Details of User Study
This section complements Section 4.1 of the main paper, providing additional details on the user studies. We involved 22 participants in these studies to evaluate our results perceptually, comparing them to baseline methods. The evaluation focused on two main aspects: font consistency (Font-Con) and style consistency (Style-Con). Style-Con was evaluated in two ways: one subtype measured the consistency between the output image and the ground truth, while the other assessed the consistency of fonts across different outputs when the same font input was used. This can be seen in Questions 1 and 2 in Figure 22. Font-Con was evaluated in a similar manner, with two subtypes addressing the same two aspects. These are represented by Question 3 of Figure 22 and Question 4 of Figure 23. Each subtype had a different number of questions: 4, 2, 3, and 2. The score for each method was determined by dividing the number of votes it received by the total number of votes cast.
Appendix D More Results
D.1 Ablation on ETC-Tokens
This section supplements Section 4.3 of the main paper, focusing on demonstrating the effectiveness of the proposed Enclosing Typography Control (ETC)-tokens for targeted word-level typographic attributes. For instance, to bold the word ”robot” in the phrase “i am not a robot”, we explore three settings: 1) Non-Token: Using a instruction prompt instead of adding modifier tokens, such as “Make ‘robot’ bold”. 2) Single-Token: Drawing on prior research [16, 6], we trained our model to use a single token, placing the modifier token before “robot”. 3) Our ETC-Token. The quantitative comparative results are presented in Table 5.
[…]
Non-Token 71.00 25.00
Single-Token 72.88 32.00
ETC-Token(Ours) 81.43 55.00
Table 5: Ablation studies of ETC-Token on basic text rendering.
D.2 Typographic Controls in STR
We found that the typography controls acquired from Basic Text Rendering (BTR) can be partially transferred to other text rendering tasks. The model’s capacity to learn typography attributes from simple text images shows considerable promise for generalization and adaptability in various domains. Consequently, this enables the application of typographic controls, as depicted in Figure 17, and font selection, as displayed in Figure 18, in Scene Text Rendering (STR).
D.3 More Qualitative Results of ATR
This section supplements Section 4.1 of the main paper, offering a qualitative comparison of our method with Glyph-ByT5 [20] and Textdiffuser-2 [9] on the ATR-bench dataset, as depicted in Figure 19. Additionally, our extended qualitative results on the ATR-bench dataset, including single-word and multi-word examples, are presented in Figure 20.
D.4 Semantic Confusion
The term “semantic confusion” in the main paper refers to instances where text rendering incorrectly generates visual objects based on the semantic meaning of the text, rather than just producing the text itself. For example, as shown in Figure 16, our intention was to render only the artistic text “Octopus”, “MOON”, and “CANDLE” in the left three images. However, the images inadvertently include the corresponding objects for these words. Similarly, in the right three images, which are supposed to display text on the scene, the text is absent, and only the specific objects associated with the semantic meaning of text are present.
Appendix E Details about Word-Acc
Current open-source OCR tools lack the capability to recognize word-level attributes such as bold, italic, and underline. To address this limitation, we employ GPT-4o [26] to evaluate the accuracy of word-level attributes (Word-Acc). We have designed a structured prompt, supplemented with example cases, to improve GPT-4o’s precision in predicting these attributes. Figure 21 illustrates a dialogue record that showcases GPT-4o’s strong context comprehension and logical reasoning abilities.
Figure 18: Examples of Font Selection in Scene Text Rendering.
Figure 19: Results of Glyph-ByT5 [20] and Textdiffuser-2 [9] on ATR-bench.
Figure 20: More Qualitative Results of Artistic Text Rendering.
Figure 21: Example of using GPT-4o to evaluate word-level attribute accuracy (Word-Acc).
Figure 22: Examples of questionnaire to evaluate the Style-Con and Font-Con.
Figure 23: Examples of questionnaire to evaluate the Font-Con.
Image Text
SC-general, Row 1, Col 1 A photorealistic image of multiple shopping bags in a boho style, fresh and inviting. The bags are in various sizes and patterns, including floral designs, abstract prints, and earthy tones. They have rope handles and are arranged against a soft, neutral background. The overall vibe is natural, stylish, and vibrant.
SC-general, Row 1, Col 2 Dark blue purple red abstract background for design. Painted rough paper. Bright colors include magenta and fuchsia. Smudge, stain, and blot effects are photo-realistic with ultra sharp focus and ultra detailed focus. The image has high coherence and minimalistic style with intricate and hyper realistic details. Beautifully color graded with modern and cinematic light. Captured with a Phase One XF IQ4 camera, 200 Mega Pixels, it features insane detailing and depth of field. The textures give a feeling of depth and richness, enhancing the overall beauty of the composition. The editorial photography and photoshoot elements are evident in the detailed and professional capture.
SC-general, Row 1, Col 3 White and grayish Persian cat with fluffy fur, vibrant green eyes, not a flat nose, has a distinct stop, looking directly into the camera, soft dramatic lighting, cinematic style, slightly backlit.
SC-general, Row 1, Col 4 an alien cyborg with eyes and oozing in the woods, in the style of rendered in cinema4d, undefined anatomy, tangled nests, dark white and crimson, eerily realistic, soft sculptures, made of mist.
SC-general, Row 1, Col 5 A luminous figure draped in glowing robes holds a radiant orb of light with plants and leaves on their shoulders, resembling the Keeper of the Light, Dota 2, in an enchanting, mystical forest ambiance.
SC-artext, Row 1, Col 1 The image presents a simple yet striking visual. Dominating the frame is the word “Average”, spelled out in capital letters. Each letter is identical in size and color, creating a sense of uniformity and balance. The letters are not solid but rather composed of small bumps, giving them a textured appearance that stands out against the stark white background. The word “Average” is centrally positioned, drawing the viewer’s attention immediately to it. Despite the simplicity of the elements involved, the image conveys a clear message: the word “Average”. The absence of any other elements or distractions underscores this message, making it the sole focus of the viewer’s attention.
SC-artext, Row 1, Col 2 The image presents a 3D rendering of the word “Average”. The word is written in a cursive font and is colored in a vibrant shade of blue. It’s slightly tilted to the right, adding a dynamic touch to the overall composition. Each letter is slightly larger than the last, creating a cascading effect that leads the viewer’s eye down the word. The background is a stark white, which contrasts sharply with the blue of the word, making it stand out prominently. The image does not contain any other objects or text, and the focus is solely on the word ”Average”. The simplicity of the image allows the viewer to clearly see and understand the meaning of the word.
SC-artext, Row 1, Col 3 The image presents a 3D rendering of the word “Average”. The word is written in a bold, sans-serif font and is colored in a vibrant shade of red. The letters are slightly tilted to the right, adding a dynamic touch to the overall composition. Each letter is enveloped in a ring of fire, with the letters “A”, “V”, and “R” being particularly noticeable due to their larger size. The background is a stark white, which contrasts sharply with the fiery red of the word, making it stand out prominently. The image does not contain any other discernible objects or text. The focus is solely on the word “Average” and its fiery presentation.
SC-artext, Row 1, Col 4 The image presents a 3D rendering of the word “Average” in a vibrant shade of green. The letters are intricately crafted from grass, giving them a natural and organic feel. Each letter is adorned with small white flowers, adding a touch of whimsy to the overall design. The letters are arranged in a staggered formation, creating a sense of depth and dimension. The word “Average” stands out prominently against the stark white background, making it the focal point of the image. The image does not contain any discernible text apart from the word “Average”.
SC-artext, Row 1, Col 5 The image presents a vibrant display of the word “Average” in a cursive font. The letters are filled with splashes of paint in a rainbow of colors, transitioning from red to orange, then to yellow, green, blue, and finally to purple. Each letter is slightly tilted, adding a dynamic feel to the overall composition. The background is a stark white, which contrasts with the colorful text and allows it to stand out prominently. The word “Average” is the only text present in the image. The relative positions of the letters suggest they are stacked on top of each other, further enhancing the visual impact of the image.
Table 6: Examples of texts in SC-general and SC-artext. Textual description of the first row in Figure 14 and 15.